Tell how many electrons there can be, in a given atom, with the following sets of quantum numbers.

1. n=3

2. n=2, l=1

3. n=4, l=2, ml =0

4. n=3, l=2, ml =1, ms =+1/2

5. n=7


1
Expert's answer
2021-11-18T02:48:02-0500

n=4 ; it can have maximum of 2n^2 electrons . = 2×4^2 = 32


m=0 : means s orbital so it can have 2 electrons


n=3,l=4; this quantum number is impossible because when n=3 , l=0,1,2 only. so # of electrons =0


l=1 ; means p subshell . It can have 6 electrons